Find the laplace transform of sinat where a is a constant


1
Expert's answer
2021-06-21T16:34:06-0400

"\begin{aligned}
\mathcal{L}[f(t)] &=F(s)=\int_{0}^{\infty} f(t) e^{-s t} d t=\int_{0}^{\infty} t^{2} \sin (a t) e^{-s t} d t \\
&=\lim _{A \rightarrow \infty} \int_{0}^{A} t^{2} \sin (a t) e^{-s t} d t\\
&=\left[\begin{array}{cc}
u=t & d v=t \sin (a t) e^{-s t} d t \\
d u=d t & v=\int t \sin (a t) e^{-s t} d t
\end{array}\right]--(*)
\end{aligned}"


"\begin{aligned}
v &=\int t \sin (a t) e^{-s t} d t=\frac{1}{2 i}\left[\int t e^{(i a-s) t} d t-\int t e^{-(i a+s) t} d t\right] \\
&=\frac{1}{2 i}\left[t \frac{e^{(i a-s) t}}{i a-s}-\int \frac{e^{(i a-s) t}}{i a-s} d t+t \frac{e^{-(i a+s) t}}{i a+s}-\int \frac{e^{-(i a+s) t}}{i a+s} d t\right] \\
&=\frac{1}{2 i}\left[t \frac{e^{(i a-s) t}}{i a-s}-\frac{e^{(i a-s) t}}{(i a-s)^{2}}+t \frac{e^{-(i a+s) t}}{i a+s}+\frac{e^{-(i a+s) t}}{(i a+s)^{2}}\right]
\end{aligned} --(i)"


Integrating v, we obtain;

"\int v d t=\frac{1}{2 i}\left[\int t \frac{e^{(i a-s) t}}{i a-s} d t-\frac{e^{(i a-s) t}}{(i a-s)^{3}}+\int t \frac{e^{-(i a+s) t}}{i a+s} d t-\frac{e^{-(i a+s) t}}{(i a+s)^{3}}\right]--(ii)"

"\begin{array}{l}
=\frac{1}{2 i}\left\{\frac{1}{i a-s}\left[t \frac{e^{(i a-s) t}}{i a-s}-\frac{e^{(i a-s) t}}{(i a-s)^{2}}\right]-\frac{e^{(i a-s) t}}{(i a-s)^{3}}\right. \\
\left.-\frac{1}{i a+s}\left[t \frac{e^{-(i a+s) t}}{i a+s}+\frac{e^{-(i a+s) t}}{(i a+s)^{2}}\right]-\frac{e^{-(i a+s) t}}{(i a+s)^{3}}\right\} \\
=\frac{1}{2 i}\left[t \frac{e^{(i a-s) t}}{(i a-s)^{2}}-2 \frac{e^{(i a-s) t}}{(i a-s)^{3}}-t \frac{e^{-(i a+s) t}}{(i a+s)^{2}}-2 \frac{e^{-(i a+s) t}}{(i a+s)^{3}}\right]
\end{array}"



Returning back to the start and using 1 and 2, we have

"\begin{aligned}& \lim _{A \rightarrow \infty}\left(\left.u v\right|_{0} ^{A}-\int_{0}^{A} v d u\right) \\
=& \lim _{A \rightarrow \infty}\left\{\left.\frac{t}{2 i}\left[t \frac{e^{(i a-s) t}}{i a-s}-\frac{e^{(i a-s) t}}{(i a-s)^{2}}+t \frac{e^{-(i a+s) t}}{i a+s}+\frac{e^{-(i a+s) t}}{(i a+s)^{2}}\right]\right|_{0} ^{A}\right.\\
&\left.-\left.\frac{1}{2 i}\left[t \frac{e^{(i a-s) t}}{(i a-s)^{2}}-2 \frac{e^{(i a-s) t}}{(i a-s)^{3}}-t \frac{e^{-(i a+s) t}}{(i a+s)^{2}}-2 \frac{e^{-(i a+s) t}}{(i a+s)^{3}}\right]\right|_{0} ^{A}\right\} \\
=& \lim _{A \rightarrow \infty}\left\{\frac{A}{2 i}\left[\frac{A e^{(i a-s) A}}{i a-s}-\frac{e^{(i a-s) A}}{(i a-s)^{2}}+\frac{A e^{-(i a+s) A}}{i a+s}+\frac{e^{-(i a+s) A}}{(i a+s)^{2}}\right]\right.\\
&-\frac{1}{2 i}\left[\frac{A e^{(i a-s) A}}{(i a-s)^{2}}-2 \frac{e^{(i a-s) A}}{(i a-s)^{3}}-\frac{A e^{-(i a+s) A}}{(i a+s)^{2}}-2 \frac{e^{-(i a+s) A}}{(i a+s)^{3}}\right] \\
&\left.+\frac{1}{2 i}\left[-2 \frac{1}{(i a-s)^{3}}-2 \frac{1}{(i a+s)^{3}}\right]\right\}
\end{aligned}"


"\begin{aligned}
F(s) &=\frac{-2}{2 i}\left[\frac{1}{(i a-s)^{3}}+\frac{1}{(i a+s)^{3}}\right]=\frac{-1}{i} \frac{(i a+s)^{3}+(i a-s)^{3}}{[(i a-s)(i a+s)]^{3}} \\
&=\frac{-1}{i} \frac{-i a^{3}-3 a^{2} s+3 i a s^{2}+s^{3}+\left(-i a^{3}+3 a^{2} s+3 i a s^{2}-s^{3}\right)}{\left(-a^{2}-s^{2}\right)^{3}} \\
&=\frac{-1}{i} \frac{-2 i a^{3}+6 i a s^{2}}{-\left(a^{2}+s^{2}\right)^{3}}=\frac{2 a\left(3 s^{2}-a^{2}\right)}{\left(a^{2}+s^{2}\right)^{3}}
\end{aligned}"